1. In a certain examination, 72 candidates offered Maths, 64 offered English, 62 offered French, 18 offered both Maths and English, 24 offered Maths and French, and 20 offered English and French. While 8 candidates offered all the three subjects. How many candidates were there for the examination?
1
Expert's answer
2020-05-06T20:09:01-0400

All number of people = Maths + English + French - (Maths and English) - (Maths and French) - (English and French) + (All three subject) = 72 + 64 + 62 - 18 - 24 - 20 + 8 = 144

Answer: 144